OpenStudy (anonymous):

Simplify 1 over quantity x-3 plus 4 over x all over 4 over x-1 over quantity x-3

OpenStudy (anonymous):

no adding... you need to find the common denominator i.e. the GCF of the denominator

OpenStudy (anonymous):

you just coppyed that from another answer ive seen that

OpenStudy (whpalmer4):

\[\frac{(\frac{1}{x-3}+\frac{4}{x})}{(\frac{(\frac{4}{x-1})}{x-3})}\]is that right?

OpenStudy (anonymous):